# TV homogenization inequalities

### v2

@@ Abstract
1 Introduction
For $\textup{{\bf p}}=(p_{1},\dots,p_{n})\in[0,1]^{n}$, the Poisson–binomial law, introduced by [14] [^1] See [15] for a modern English translation. , is the pushforward of the product measure $\operatorname{Ber}(\textup{{\bf p}}):=\operatorname{Ber}(p_{1})\otimes\cdots\otimes\operatorname{Ber}(p_{n})$ under the sum map $x\mapsto\sum_{i}x_{i}$. Equivalently, if $X_{i}\sim\operatorname{Ber}(p_{i})$ are independent, then $S_{\textup{{\bf p}}}:=\sum_{i=1}^{n}X_{i}$ is distributed according to this law. The sum map is easily seen to be a Markov kernel, and hence the Data Processing Inequality [16, Theorem 7.4] applies: $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, where $\operatorname{TV}$ is the total variation distance. Intuitively, since information is lost under the sum map, the distributions can only become closer to each other. In this paper, we investigate the behavior of TV under the sum map at the parameter rather than the observation level. The homogenization map replaces each $p_{i}$ with $\bar{p}=\frac{1}{n}\sum_{i}p_{i}$, transforming $\operatorname{Ber}(\textup{{\bf p}})$ into $\operatorname{Ber}(\bar{p})^{\otimes n}$. Since this operation also erases information, we might likewise expect it to decrease the TV distance. One immediate obstruction is that the homogenization map cannot be realized by any Markov kernel. Indeed, suppose to the contrary that for $n\geq 2$, some Markov kernel $K$ maps $X\sim\operatorname{Ber}(\textup{{\bf p}})$ to $Y\sim\operatorname{Ber}(\bar{p})^{\otimes n}$. Applying the sum map to $Y$ yields $Z\sim\operatorname{Bin}(n,\bar{p})$; thus we have some Markov kernel $\tilde{K}:\operatorname{Ber}(\textup{{\bf p}})\mapsto\operatorname{Bin}(n,\bar{p})$ for all $\textup{{\bf p}}\in[0,1]^{n}$. Consider a p of the form $\textup{{\bf p}}=(t,0,0,\ldots,0)\in[0,1]^{n}$. Then the law of $X$ is affine in $t$ (it is a $t$-convex combination of two atoms), as is the pushforward measure under $\tilde{K}$. However, $\mathbb{P}(Z=2)=\binom{n}{2}\left(\frac{t}{n}\right)^{2}\left(1-\frac{t}{n}\right)^{n-2}$, which is not affine in tt. [^2] For the special case of $\textup{{\bf p}},\textup{{\bf q}}$ that are symmetric about $\frac{1}{2}$, meaning that $\textup{{\bf q}}=\textup{{\bf 1}}-\textup{{\bf p}}$, it turns out that the homogenization operator can be realized as a Markov kernel; see Appendix.
@@ 1 Introduction
$\Delta:=\sum_{i=1}^{n}|p_{i}-q_{i}|,\quad\sigma_{\textup{{\bf p}}}^{2}:=\operatorname{Var}(S_{\textup{{\bf p}}})=\sum_{i=1}^{n}p_{i}(1-p_{i}),\quad\Phi(\textup{{\bf p}},\textup{{\bf q}}):=\min\!\left(1,\frac{\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}\right).$ (1)
@@
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp\Phi(\textup{{\bf p}},\textup{{\bf q}}),\qquad\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n},\textup{{\bf p}}\geq\textup{{\bf q}}.$ (2)
For general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, only the upper bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\lesssim\Phi(\textup{{\bf p}},\textup{{\bf q}})$ holds, which can be sharpened to $\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}}))$ by symmetry. Unlike $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$, whose exact computation is infeasible (see below), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ is computable exactly in time $O(n^{2})$ via a simple recurrence [20, Section 5]. Still, one should not expect any analytically tractable analogue of (2) for general $\textup{{\bf p}},\textup{{\bf q}}$ due to cancellations and convolutional smoothing. As an application of these results, we prove the TV homogenization inequality for all $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$:
$\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\operatorname{TV}(\operatorname{Ber}(\bar{p})^{\otimes n},\operatorname{Ber}(\bar{q})^{\otimes n})=\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})).$ (3)
Notation.
We write $n]:=\left\{1,\ldots,n\right\}$. For two distributions $P,Q$ on a finite set $\Omega$, their total variation distance is defined by
$\operatorname{TV}(P,Q)$ $=$ $\frac{1}{2}\sum_{\omega\in\Omega}|P(\omega)-Q(\omega)|.$ (4)
@@
Main results.
Our first result implies the upper bound in (2). Let $\eta_{\mathrm{BCV}}$ denote the sharp universal constant from [1, Theorem 1], and set
$C_{\mathrm{BCV}}:=\sqrt{\frac{5}{4}+\eta_{\mathrm{BCV}}^{2}}.$
Numerically (see [1, Remark 1]), $\eta_{\mathrm{BCV}}\approx 0.4688223555$ and hence $C_{\mathrm{BCV}}\approx 1.2123507747$.
@@
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $2C_{\mathrm{BCV}}\,\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}})).$
@@
The assumption $\textup{{\bf p}}\geq\textup{{\bf q}}$ is essential for the lower bound, as illustrated by the example $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, for even $n$, given by
$\textup{{\bf p}}=\Bigl(\tfrac{1}{2},\tfrac{1}{2},\dots,\tfrac{1}{2}\Bigr)$, $\textup{{\bf q}}=\Bigl(\tfrac{1}{2}+\delta,\dots,\tfrac{1}{2}+\delta,\ \tfrac{1}{2}-\delta,\dots,\tfrac{1}{2}-\delta\Bigr),$
with $n/2$ copies of $\tfrac{1}{2}+\delta$ followed by $n/2$ copies of $\tfrac{1}{2}-\delta$.
The choice
$\delta=1/\sqrt{n}$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})=1$, but $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp 1/n$. In general, for any $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and any partition $I,J$ of $[n]$, the trivial estimate via the partitioned pairs $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\leq\operatorname{TV}(S_{\textup{{\bf p}}_{I}},S_{\textup{{\bf q}}_{I}})+\operatorname{TV}(S_{\textup{{\bf p}}_{J}},S_{\textup{{\bf q}}_{J}})$ holds. However, this method cannot yield a matching lower bound, as illustrated by the example $a,b=\tfrac{1}{2}\pm\varepsilon$, $\textup{{\bf p}}=(a,a,b,b)$, $\textup{{\bf q}}=(a,b,a,b)$ with the partition $I=\left\{1,2\right\}$, $J=\left\{3,4\right\}$. Although for general $\textup{{\bf p}},\textup{{\bf q}}$, the approach of Theorem 1.2 does not lower-bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, it does yield a lower bound for the larger quantity $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$:
@@
where $c\geq\frac{1}{72C_{\mathrm{BCV}}}\approx 0.0115$ is a universal constant.
Remark.
As the remark preceding Theorem 2.1 indicates, a sharpened constant $c\geq\frac{1}{36C_{\mathrm{BCV}}}$ is straightforward to obtain via analogous arguments. The choice $\textup{{\bf p}}=(1-2\varepsilon,\tfrac{1}{2})$, $\textup{{\bf q}}=(1,\tfrac{1}{2}+\varepsilon)$, $\varepsilon\to 0$ demonstrates that it is possible for homogenization to (slightly) increase the TV distance, and also that $8/9$ is an upper bound on $c$. We conjecture that this value is in fact optimal, see below.
@@
$\geq$ $\frac{1}{2C_{\mathrm{BCV}}}\left(\operatorname{TV}(\operatorname{Bin}(|I|,\bar{p}_{I}),\operatorname{Bin}(|I|,\bar{q}_{I}))+\operatorname{TV}(\operatorname{Bin}(|J|,\bar{p}_{J}),\operatorname{Bin}(|J|,\bar{q}_{J}))\right)$
$\geq$ $\frac{1}{4C_{\mathrm{BCV}}}\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
@@
Although the homogenization inequality relies on novel structural insights into the Poisson binomial, there are compelling reasons to believe that a great deal more structure remains to be uncovered, currently out of reach. Indeed, extensive numerical experiments suggest that the correct constant in Theorem 1.5 should be $c=8/9$ and also that the bound in Lemma 1.4 can be sharpened to $\delta_{N}\leq\delta_{I}+\delta_{J}-\delta_{I}\delta_{J}$ (the latter, in particular, appears deceptively simple, since only binomials are involved). Our present methods do not seem to provide any pathway to these conjecturally optimal bounds, which will have to await further structural advances.
Related work.
@@
This paper follows a program initiated by [9] to provide simple, analytically tractable upper and lower estimates on $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ in terms of the $\textup{{\bf p}},\textup{{\bf q}}$; the latter’s contribution consisted of sharpening of the trivial lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{\infty}$ to $\gtrsim\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}$. We note that generally, this result is incomparable with the lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\Phi(\textup{{\bf p}},\textup{{\bf q}})$ implied by Theorem 1.2 for $\textup{{\bf p}}\geq\textup{{\bf q}}$. Indeed, the choice $q_{i}\equiv\frac{1}{2}$ for $i\in[n]$ and $p_{i}=1/2$, $i>1$ and $p_{1}=\frac{1}{2}+\varepsilon$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp\varepsilon/\sqrt{n}$ while $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}=\varepsilon$. For the other direction, the choice $\textup{{\bf q}}\equiv 0$ and $\textup{{\bf p}}\equiv 1/n$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp 1$ and $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|=n^{-1/2}$. This program is continued in [10], where an analytical closed-form $O(\sqrt{\log n})$-factor approximation to $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ is obtained.
@@
Homogenization inequalities appear not to have been widely studied; one classic result due to Hoeffding [8] is that homogenization under a fixed-mean constraint maximizes $\mathbb{E}g(S_{\textup{{\bf p}}})$ for any convex $g$. The study of $S_{\textup{{\bf p}}}$ is amply motivated by numerous applications and a venerable line of work has considered approximating this distribution by simpler ones [20]; perhaps most famous is Le Cam’s inequality $\operatorname{TV}(S_{\textup{{\bf p}}},\mathrm{Poi}(\sum p_{i}))\leq\sum{p_{i}^{2}}$ [11]. Note, however, that the latter yields an additive, rather than multiplicative approximation to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$. Binomial approximation bounds in strong metrics (including $\operatorname{TV}$) appear in [5] and refinements based on orthogonal-polynomial expansions and asymptotics in [17], then generalized to arbitrary laws [18] and also gave an additive homogenization-type estimate. Later, [19] obtained TV bounds for approximating general convolutions by compound Poisson laws using Kerstan’s method together with novel “smoothness inequalities”. A delicate structural result due to [4] (also ultimately yielding an additive estimate) shows that every Poisson binomial admits an $\varepsilon$-approximation in TV by a distribution of one of two canonical “compressed” types: either a sparse $S_{\textup{{\bf p}}}$ with only $O(1/\varepsilon^{3})$ nontrivial summands, or a near-binomial form. Of particular relevance to this work is [1], whose structural results we build upon in Theorem 1.1. More generally, $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ admits classic approximations in terms of more analytically tractable proxies such as KL-divergence and the Hellinger distance; their properties and limitations are discussed in [9]. Regarding the algorithmic aspect, [3] showed that computing $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ exactly for general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ is hard in the #\#P sense. An efficient randomized algorithm guaranteeing a $1\pm\varepsilon$ multiplicative approximation with confidence $1-\delta$, in time $O(\frac{n^{2}}{\varepsilon^{2}}\log\frac{1}{\delta})$ was discovered by [6], and later derandomized by [7].
2 Proofs
2.1 Proof of Theorem 1.1

### v3

@@ Abstract.
Key words and phrases:
total variation; homogenization; Poisson binomial; Bernoulli product measures
2020 Mathematics Subject Classification:
60E15, 60F05
1. Introduction
For $\textup{{\bf p}}=(p_{1},\dots,p_{n})\in[0,1]^{n}$, the Poisson–binomial law, introduced by [16] [^1] See [17] for a modern English translation. , is the pushforward of the product measure $\operatorname{Ber}(\textup{{\bf p}}):=\operatorname{Ber}(p_{1})\otimes\cdots\otimes\operatorname{Ber}(p_{n})$ under the sum map $x\mapsto\sum_{i}x_{i}$. Equivalently, if $X_{i}\sim\operatorname{Ber}(p_{i})$ are independent, then $S_{\textup{{\bf p}}}:=\sum_{i=1}^{n}X_{i}$ is distributed according to this law. The sum map is easily seen to be a Markov kernel, and hence the Data Processing Inequality [18, Theorem 7.4] applies: $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, where $\operatorname{TV}$ is the total variation distance. Intuitively, since information is lost under the sum map, the distributions can only become closer to each other. In this paper, we investigate the behavior of TV under the sum map at the parameter rather than the observation level. The homogenization map replaces each $p_{i}$ with $\bar{p}=\frac{1}{n}\sum_{i}p_{i}$, transforming $\operatorname{Ber}(\textup{{\bf p}})$ into $\operatorname{Ber}(\bar{p})^{\otimes n}$. Since this operation also erases information, we might likewise expect it to decrease the TV distance. One immediate obstruction is that the homogenization map cannot be realized by any Markov kernel. Indeed, suppose to the contrary that for $n\geq 2$, some Markov kernel $K$ maps $X\sim\operatorname{Ber}(\textup{{\bf p}})$ to $Y\sim\operatorname{Ber}(\bar{p})^{\otimes n}$. Applying the sum map to $Y$ yields $Z\sim\operatorname{Bin}(n,\bar{p})$; thus we have some Markov kernel $\tilde{K}:\operatorname{Ber}(\textup{{\bf p}})\mapsto\operatorname{Bin}(n,\bar{p})$ for all $\textup{{\bf p}}\in[0,1]^{n}$. Consider a p of the form $\textup{{\bf p}}=(t,0,0,\ldots,0)\in[0,1]^{n}$. Then the law of $X$ is affine in $t$ (it is a $t$-convex combination of two atoms), as is the pushforward measure under $\tilde{K}$. However, $\mathbb{P}(Z=2)=\binom{n}{2}\left(\frac{t}{n}\right)^{2}\left(1-\frac{t}{n}\right)^{n-2}$, which is not affine in tt. [^2] For the special case of complementary $\textup{{\bf p}},\textup{{\bf q}}$, meaning that $\textup{{\bf q}}=\textup{{\bf 1}}-\textup{{\bf p}}$, it turns out that the homogenization operator can be realized as a Markov kernel; see Appendix.
@@ 1. Introduction
(1) $\Delta:=\sum_{i=1}^{n}|p_{i}-q_{i}|,\quad\sigma_{\textup{{\bf p}}}^{2}:=\operatorname{Var}(S_{\textup{{\bf p}}})=\sum_{i=1}^{n}p_{i}(1-p_{i}),\quad\Phi(\textup{{\bf p}},\textup{{\bf q}}):=\min\!\left(1,\frac{\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}}\right).$
@@
(2) $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp\Phi(\textup{{\bf p}},\textup{{\bf q}}),\qquad\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n},\textup{{\bf p}}\geq\textup{{\bf q}}.$
For general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, only the upper bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\lesssim\Phi(\textup{{\bf p}},\textup{{\bf q}})$ holds, which can be sharpened to $\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}}))$ by symmetry. Unlike $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$, whose exact computation is infeasible (see below), $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ is computable exactly in time $O(n^{2})$ via a simple recurrence [22, Section 5]. Still, one should not expect any analytically tractable analogue of (2) for general $\textup{{\bf p}},\textup{{\bf q}}$ due to cancellations and convolutional smoothing. As an application of these results, we prove the TV homogenization inequality for all $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$:
(3) $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\operatorname{TV}(\operatorname{Ber}(\bar{p})^{\otimes n},\operatorname{Ber}(\bar{q})^{\otimes n})=\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})).$
Notation. We write $n]:=\left\{1,\ldots,n\right\}$. For two distributions $P,Q$ on a finite set $\Omega$, their total variation distance is defined by
(4) $\operatorname{TV}(P,Q)$ $=$ $\frac{1}{2}\sum_{\omega\in\Omega}|P(\omega)-Q(\omega)|.$
@@
Main results. Our first result implies the upper bound in (2). Let $\eta_{0}$ denote the sharp universal constant from [2, Theorem 1], and set
$C_{0}:=\sqrt{\frac{5}{4}+\eta_{0}^{2}}.$
Numerically (see [2, Remark 1]), $\eta_{0}\approx 0.4688223555$ and hence $C_{0}\approx 1.2123507747$.
@@
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $2C_{0}\,\min(\Phi(\textup{{\bf p}},\textup{{\bf q}}),\Phi(\textup{{\bf q}},\textup{{\bf p}})).$
L. Mattner [14] has brought to our attention the result of [23], which upper bounds $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ by
@@
The assumption $\textup{{\bf p}}\geq\textup{{\bf q}}$ is essential for the lower bound, as illustrated by the example $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$, for even $n$, given by $\textup{{\bf p}}=\Bigl(\tfrac{1}{2},\tfrac{1}{2},\dots,\tfrac{1}{2}\Bigr)$, $\textup{{\bf q}}=\Bigl(\tfrac{1}{2}+\delta,\dots,\tfrac{1}{2}+\delta,\ \tfrac{1}{2}-\delta,\dots,\tfrac{1}{2}-\delta\Bigr),$ with $n/2$ copies of $\tfrac{1}{2}+\delta$ followed by $n/2$ copies of $\tfrac{1}{2}-\delta$. The choice $\delta=1/\sqrt{n}$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})=1$, but $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\asymp 1/n$. In general, for any $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ and any partition $I,J$ of $[n]$, the trivial estimate via the partitioned pairs $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\leq\operatorname{TV}(S_{\textup{{\bf p}}_{I}},S_{\textup{{\bf q}}_{I}})+\operatorname{TV}(S_{\textup{{\bf p}}_{J}},S_{\textup{{\bf q}}_{J}})$ holds. However, this method cannot yield a matching lower bound, as illustrated by the example $a,b=\tfrac{1}{2}\pm\varepsilon$, $\textup{{\bf p}}=(a,a,b,b)$, $\textup{{\bf q}}=(a,b,a,b)$ with the partition $I=\left\{1,2\right\}$, $J=\left\{3,4\right\}$. Although for general $\textup{{\bf p}},\textup{{\bf q}}$, the approach of Theorem 1.2 does not lower-bound $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$, it does yield a lower bound for the larger quantity $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$:
@@
where $c\geq\frac{1}{72C_{0}}\approx 0.0115$ is a universal constant.
Remark. As the remark preceding Theorem 2.1 indicates, a sharpened constant $c\geq\frac{1}{36C_{0}}$ is straightforward to obtain via analogous arguments. The choice $\textup{{\bf p}}=(1-2\varepsilon,\tfrac{1}{2})$, $\textup{{\bf q}}=(1,\tfrac{1}{2}+\varepsilon)$, $\varepsilon\to 0$ demonstrates that it is possible for homogenization to (slightly) increase the TV distance, and also that $8/9$ is an upper bound on $c$. We conjecture that this value is in fact optimal, see below.
@@
$\geq$ $\frac{1}{2C_{0}}\left(\operatorname{TV}(\operatorname{Bin}(|I|,\bar{p}_{I}),\operatorname{Bin}(|I|,\bar{q}_{I}))+\operatorname{TV}(\operatorname{Bin}(|J|,\bar{p}_{J}),\operatorname{Bin}(|J|,\bar{q}_{J}))\right)$
$\geq$ $\frac{1}{4C_{0}}\operatorname{TV}(\operatorname{Bin}(n,\bar{p}),\operatorname{Bin}(n,\bar{q})),$
@@
Remark and open problems. Although the homogenization inequality relies on novel structural insights into the Poisson binomial, there are compelling reasons to believe that a great deal more structure remains to be uncovered, currently out of reach. Indeed, extensive numerical experiments suggest that the correct constant in Theorem 1.5 should be $c=8/9$ and also that the bound in Lemma 1.4 can be sharpened to $\delta_{N}\leq\delta_{I}+\delta_{J}-\delta_{I}\delta_{J}$ (the latter, in particular, appears deceptively simple, since only binomials are involved). Our present methods do not seem to provide any pathway to these conjecturally optimal bounds, which will have to await further structural advances.
@@
Related work. This paper follows a program initiated by [11] to provide simple, analytically tractable upper and lower estimates on $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ in terms of the $\textup{{\bf p}},\textup{{\bf q}}$; the latter’s contribution sharpened the trivial lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))\geq\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{\infty}$ to $\gtrsim\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}$. We note that generally, this result is incomparable with the lower bound $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ $\gtrsim$ $\Phi(\textup{{\bf p}},\textup{{\bf q}})$ implied by Theorem 1.2 for $\textup{{\bf p}}\geq\textup{{\bf q}}$. Indeed, the choice $q_{i}\equiv\frac{1}{2}$ for $i\in[n]$ and $p_{i}=1/2$, $i>1$ and $p_{1}=\frac{1}{2}+\varepsilon$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp\varepsilon/\sqrt{n}$ while $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|_{2}=\varepsilon$. For the other direction, the choice $\textup{{\bf q}}\equiv 0$ and $\textup{{\bf p}}\equiv 1/n$ yields $\Phi(\textup{{\bf p}},\textup{{\bf q}})\asymp 1$ and $\left\|\textup{{\bf p}}-\textup{{\bf q}}\right\|=n^{-1/2}$. This program is continued in [12], where an analytical closed-form $O(\sqrt{\log n})$-factor approximation to $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ is obtained.
@@
Homogenization inequalities appear not to have been widely studied; one classic result due to Hoeffding [10] is that homogenization under a fixed-mean constraint maximizes $\mathbb{E}g(S_{\textup{{\bf p}}})$ for any convex $g$. The study of $S_{\textup{{\bf p}}}$ is amply motivated by numerous applications and a venerable line of work has considered approximating this distribution by simpler ones [22]; perhaps most famous is Le Cam’s inequality $\operatorname{TV}(S_{\textup{{\bf p}}},\mathrm{Poi}(\sum p_{i}))\leq\sum{p_{i}^{2}}$ [13]. Note, however, that the latter yields an additive, rather than multiplicative approximation to $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$. Binomial approximation bounds in strong metrics (including $\operatorname{TV}$) appear in [7] and refinements based on orthogonal-polynomial expansions and asymptotics in [19], then generalized to arbitrary laws [20] and also gave an additive homogenization-type estimate. Later, [21] obtained TV bounds for approximating general convolutions by compound Poisson laws using Kerstan’s method together with novel “smoothness inequalities”. A delicate structural result due to [6] (also ultimately yielding an additive estimate) shows that every Poisson binomial admits an $\varepsilon$-approximation in TV by a distribution of one of two canonical “compressed” types: either a sparse $S_{\textup{{\bf p}}}$ with only $O(1/\varepsilon^{3})$ nontrivial summands, or a near-binomial form. Of particular relevance to this work is [2, 3], whose structural results we build upon in Theorem 1.1. More generally, $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ admits classic approximations in terms of more analytically tractable proxies such as KL-divergence and the Hellinger distance; their properties and limitations are discussed in [11]. Regarding the algorithmic aspect, [4] showed that computing $\operatorname{TV}(\operatorname{Ber}(\textup{{\bf p}}),\operatorname{Ber}(\textup{{\bf q}}))$ exactly for general $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ is hard in the #\#P sense. An efficient randomized algorithm guaranteeing a $1\pm\varepsilon$ multiplicative approximation with confidence $1-\delta$, in time $O(\frac{n^{2}}{\varepsilon^{2}}\log\frac{1}{\delta})$ was discovered by [8], and later derandomized by [9].
2. Proofs
2.1. Proof of Theorem 1.1